OpenStudy (anonymous):

5/x^2+x-2 - 6/x^2+2x-3 Reduce to the lowest terms, simplify.

OpenStudy (anonymous):

Please use parentheses, it really helps clarify which parts of the expression go together in fractions. I'm assuming the expression is (5 / (x^2 + x - 2)) - (6 / (x^2 + 2x - 3)).

OpenStudy (stacey):

Do you know how to factor either denominator?

OpenStudy (anonymous):

Sorry, & yeah that's how the expression is. & hm let me try I tried (x+1)(x+1)(x-1) for x^2+x-2 It doesn't look right to me.

OpenStudy (stacey):

It's not. Since we have x^2, we will only have two parenthesis. Also the third term being negative tells us that we have a plus and a minus. So we know that it factors into (x-__)(x+__)

OpenStudy (anonymous):

oh (x-1)(x+1) & then it'd be (x-2)(x+1) ?

OpenStudy (stacey):

To get a -2, we need the blanks must multiply to be 2, but add to be the middle number. For example: (x-1)(x+2)=x^2 + x - 2 because -1*2=-2 and -1 + 2 = 1

OpenStudy (anonymous):

oh yeah that makes sense. so for the other equation it has to add up to multiply to -3 but add up to 2 right

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

(x-1 ) (x+ 3)

OpenStudy (stacey):

Very good!

OpenStudy (anonymous):

Now I got to multiply the 5 by (x+3) right to make it like dem and then the 6(x+2)

OpenStudy (anonymous):

& thank you :D

OpenStudy (stacey):

You are correct. & your're very welcome.

OpenStudy (anonymous):

You're better than my teacher! haha, let me work it out and see what I get

OpenStudy (stacey):

Thanks.

OpenStudy (anonymous):

on top :5x+15-6x+12 bottom (x-1)(x+2)(x+3)

OpenStudy (anonymous):

-x+27 ?

OpenStudy (stacey):

The bottom is correct.

OpenStudy (stacey):

On top: (5x+15) - (6x+12) Do you see how that will change your answer a little?

OpenStudy (anonymous):

I factor?!

OpenStudy (anonymous):

5(x+3) 6(x+2)

OpenStudy (stacey):

Sorry. That was not what I was trying to say. The minus gets distributed to the 6x and the 12.

OpenStudy (anonymous):

oh wait that takes me back to the step before :/

OpenStudy (anonymous):

oh yeah

OpenStudy (stacey):

So instead of 5x + 15 - 6x +12, we should have 5x + 15 - 6x - 12.

OpenStudy (anonymous):

-x-3 ?

OpenStudy (stacey):

-x + 3

OpenStudy (anonymous):

oh cause the 15 is positive

OpenStudy (stacey):

Correct

OpenStudy (anonymous):

the x+3 doesn't cancel right, because it has a - infront

OpenStudy (stacey):

Correct. Nothing cancels.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

(x^3a+1 y^3b+4) / (y^2b-3 x^2a-1) Sorry! I really want to ace this test :(

OpenStudy (stacey):

Is 3a the first exponent?

OpenStudy (anonymous):

yes x^3a+1 the 3a+1 is the exponent

OpenStudy (stacey):

Oh. That makes more sense.

OpenStudy (stacey):

When we divide with variables that are the same, we can subtract exponents. For example, x^12/x^3 = x^(12-3) = x^9

OpenStudy (anonymous):

Oh yeah then the 3a+1/2a-1 =

OpenStudy (anonymous):

3a-2a and then 1-(-1)

OpenStudy (anonymous):

Right?

OpenStudy (stacey):

Yes. The x-exponent would be (3a+1)-(2a-1) = 3a+1-2a+1 = a+2

OpenStudy (anonymous):

okay so then 3b+4-2b-3 = b+1

OpenStudy (anonymous):

x^a+2 y^b+1 right?

OpenStudy (stacey):

Remember to distribute the minus to the -3. 4-(-3) = 4+3 = 7

OpenStudy (anonymous):

o0oo yeah, then x^a+2 y^b+7

OpenStudy (stacey):

Yes.

OpenStudy (stacey):

I am going to go. Good luck on your test.

OpenStudy (anonymous):

Thank you appreciate it!

OpenStudy (stacey):

You're very welcome.